OpenStudy (anonymous):

sec^4theta-tan^4/sec^2+tan^2theta=sec^2-tan^2theta

OpenStudy (anonymous):

\[\frac{\sec^4\theta-\tan^4\theta}{\sec^2\theta+\tan^2\theta}=\sec^2\theta-\tan^2\theta\] With the LHS, factor the numerator using the difference of squares formula: \[\sec^4\theta-\tan^4\theta=\left(\sec^2\theta+\tan^2\theta\right)\left(\sec^2\theta-\tan^2\theta\right)\] The next step should be quite obvious.